OpenStudy (anonymous):

where does 2x' come? |x^2-x'^2| = |x-x'||2x'+(x-x')|

OpenStudy (anonymous):

if a = x and b = x', this is how it comes........ ( a^2-b^2) =(a-b)(a+b) =(a-b)(2a-a+b) =(a-b)(2a+(b-a))